# Unlabeled Data or Pre-trained Model: Rethinking Semi-Supervised Learning and Pretrain-Finetuning

### v3

@@ Abstract
Semi-supervised learning (SSL) alleviates the cost of data labeling process by exploiting unlabeled data, and has achieved promising results on various tasks such as image classification. Meanwhile, the Pretrain-Finetuning paradigm has garnered significant attention in recent years, and exploiting pre-trained models could also reduce the requirement of labeled data in downstream tasks. Therefore, a question naturally occurs: When the labeled data is scarce in the target tasks, should we exploit unlabeled data or pre-trained models? To answer this question, we select pre-trained Vision-Language Models (VLMs) as representative pretrain-finetuning instances and propose Few-shot SSL—a framework that enables fair comparison between these two paradigms by controlling the amount of labeled data used. Extensive experiments across various settings demonstrate that pre-trained VLMs generally outperform SSL methods in nearly all cases, except when the data has low resolution or lacks clear semantic structure. Therefore, we encourage future SSL research to compare with pre-trained models and explore deeper integration, such as using pre-trained knowledge to enhance pseudo-labeling. To support future research, we release our unified reproduction and evaluation framework. Codes are available here.
@@ 1 Introduction
Deep neural networks have demonstrated performance comparable to that of humans in certain supervised learning tasks, such as image classification [13, 20, 28]. However, these impressive results rely heavily on large amounts of labeled training data. In many real-world applications, obtaining such labeled data often demands substantial human and financial resources [54]. To reduce the dependence on labeled data and thereby cut costs, researchers have proposed Semi-Supervised Learning (SSL)—an effective approach that leverages unlabeled data to enhance model performance.
Recently, the Pretrain-Finetuning paradigm has gained significant attention, as models pre-trained on large-scale, readily available web data exhibit a strong ability to rapidly adapt to downstream tasks through fine-tuning. In image classification, for example, the semantic alignment acquired during pre-training enables models—such as CLIP [43], a pre-trained vision-language model—to perform zero-shot classification without task-specific supervision. Furthermore, building upon the rich knowledge gained during pre-training, various fine-tuning strategies (e.g., prompt tuning) have been proposed to boost task-specific performance with minimal labeled data [24, 25, 56, 57].
@@
Both SSL and pretrain-finetuning paradigms are well-suited for downstream tasks with limited labeled data, leveraging either unlabeled data or pre-trained knowledge, respectively. This makes us come up with a question naturally:
When labeled data is scarce in the target task, should we rely on additional unlabeled data or pre-trained models? What are the respective advantages and disadvantages of each paradigm?
To answer this question, we propose Few-shot SSL, an evaluation framework designed to provide a fair comparison between SSL methods and pre-trained models under consistent conditions. This framework ensures a straightforward and intuitive comparison by aligning the number of labeled samples per class used for fine-tuning with the minimum number observed across various SSL settings. We adopt Vision-Language Models (VLMs), such as CLIP and its enhanced variants, as representative pre-trained models due to their strong generalization ability and proven effectiveness in zero- and few-shot learning across diverse domains. Both paradigms are then evaluated across four widely adopted semi-supervised settings within the unified Few-shot SSL framework, providing a comprehensive analysis of their respective advantages and limitations.
Experimental results indicate that SSL methods are well-suited for handling low-resolution datasets with limited semantic information. In contrast, pretrain-finetuning paradigm, without relying on unlabeled data, achieves competitive performance across various datasets and open-environment settings. The key takeaways under different settings are as follows:
  • Pre-trained models demonstrate strong generalization capabilities and high data efficiency, enabling them to outperform traditional SSL methods with only a small amount of labeled data. However, achieving optimal performance with pre-trained VLMs requires higher-resolution input images, typically exceeding 96$\times$96.
  • When the unlabeled data contains out-of-distribution classes, pre-trained VLMs can be leveraged to avoid unlabeled training, capitalizing on their extensive pre-trained knowledge and rapid adaptation capabilities. Additionally, pre-trained VLMs offer advantages such as robust result stability and high performance.
  • In the open-world setting, SSL methods perform better on low-resolution datasets, while pre-trained VLMs demonstrate balanced performance on both base and new classes in high-resolution datasets. Combined with our open-set analysis, we conclude that the pretrain-finetuning paradigm is better suited for scenarios where unlabeled data may contain previously unseen classes.
  • The low data requirements of pre-trained models enable them to effectively handle long-tailed distributions, offering a promising approach to addressing the inherent imbalance in many real-world datasets. Meanwhile, SSL methods remain well-suited for tasks characterized by less distinct semantic boundaries between categories, where learning directly from data representations proves advantageous.
In summary, the pretrain-finetune paradigm demonstrates strong performance across most tasks, except in scenarios with limited image resolutions or insufficient semantic information. To further advance research in related areas, we propose the following directions: (I) Incorporate comparisons with pre-trained models (e.g., VLMs) in SSL studies focused on image classification task. (II) Enhance the robustness of pre-trained models to low-resolution visual inputs, thereby improving their versatility across diverse real-world scenarios. (III) Explore methods for integrating semi-supervised learning with pre-trained model adaptation to further boost performance under limited supervision.
The remainder of this paper is structured as follows: In Section 2, we review the related work on the two paradigms, followed by a detailed explanation of their underlying mechanisms and our proposed Few-shot SSL framework in Section 3. In Section 4, we present extensive experiments to thoroughly investigate the performance of both paradigms under various settings. In Section 5, we provide an in-depth discussion on potential future directions, including the integration of these two paradigms and the impact of long-tailed data distributions on pretrain-finetuning approaches. Finally, in Section 6, we conclude our paper and offer recommendations for future research.
@@ 2 Related Work
Semi-Supervised Learning. The goal of Semi-Supervised Learning is to reduce the cost of data annotation by leveraging unlabeled data, under the assumption that both labeled and unlabeled samples belong to a predefined set of known classes. The primary objective is to assign accurate ground-truth labels to unlabeled instances. Existing SSL methods can be broadly categorized into three main approaches: Entropy Minimization [14, 15], Consistency Regularization[30, 38, 46], and Hybrid or Holistic Methods [4, 5]. Specifically, EntMin [14] promotes confident (i.e., low-entropy) predictions on unlabeled samples by introducing an additional loss term that minimizes prediction uncertainty, regardless of class. Pseudo-Labeling[32] assigns pseudo-labels to unlabeled data based on the model’s own predictions throughout training. FixMatch [48] combines consistency regularization and confidence-based filtering by enforcing consistency between weakly and strongly augmented samples, while only using high-confidence pseudo-labels for training, resulting in significant performance gains with limited labeled data.
Pretrain-Finetuning. The recent progress in pre-trained models has garnered significant attention, particularly with the emergence of Large Language Models like GPT [1] and Vision Language Models like CLIP [43]. These models, pre-trained on vast amounts of data, exhibit remarkable adaptability to downstream tasks, leading to the widespread adoption of the pretrain-finetuning paradigm. In image classification tasks, VLMs [2, 43, 55] are pre-trained on large-scale text-image pairs, leveraging shared embeddings of textual and visual representations to enable zero-shot classification. While these pre-trained VLMs have captured broad, generalized representations of images and text, efficiently adapting them to downstream tasks remains a significant challenge. Inspired by the success of prompt tuning in NLP [21, 33], many researchers have proposed to adapt VLMs by learning the prompts in end-to-end training. As the pioneering work, CoOp [57] for the first time introduces learnable prompts to transfer the task-specific knowledge to VLMs. Building on this idea, MaPLe [24] and PromptSRC [25] extend prompt tuning to both visual and textual encoders, enabling joint vision-language adaptation. In Few-shot SSL, we select CoOp and PromptSRC as the primary baselines.
3 Preliminary
In this section, we first present an overview of the working principles behind SSL and the pretrain–finetuning paradigms, as illustrated in Figure 1, along with their respective strategies for adapting to downstream tasks. We then describe the four semi-supervised settings used in our experiments, both in their original form and as adapted within our Few-shot SSL framework.
3.1 Semi-Supervised Learning
SSL leverages both a small set of labeled examples $\mathcal{D}_{l}=\{(x_{i},y_{i})\}_{i=1}^{N_{l}}$ and a larger pool of unlabeled examples $\mathcal{D}_{u}=\{x_{j}\}_{j=1}^{N_{u}}$ during training stage (see Figure 1 (a)). Let $f_{\theta}$ denote the model with parameters $\theta$. SSL jointly minimizes a supervised loss on $\mathcal{D}_{l}$ and an unsupervised loss on $\mathcal{D}_{u}$:
$\mathcal{L}(\theta)=\frac{1}{N_{l}}\sum_{i=1}^{N_{l}}\ell\bigl{(}f_{\theta}(x_{i}),y_{i}\bigr{)}\;+\;\lambda\frac{1}{N_{u}}\sum_{j=1}^{N_{u}}\mathcal{L}_{u}\bigl{(}f_{\theta},x_{j}\bigr{)},$ (1)
Here, $\ell(\cdot,\cdot)$ denotes the cross-entropy loss. The first term represents the supervised loss, while the second term captures the unsupervised loss, with $\lambda>0$ serving as a balancing factor between them. Although SSL reduces the dependence on labeled data, this advantage often comes with significant computational overhead during training, owing to reliance on large-scale use of unlabeled data.
3.2 Pretrain-Fintuning
Pre-trained models typically involve two stages before downstream evaluation (see Figure 1(b)). Taking CLIP [43] as an example, it employs an image encoder $f$ and a text encoder $g$ to extract feature representations $f(I)$ and $g(T)$ from an input image $I$ and its corresponding text caption $T$, respectively. During pre-training, the model minimizes a contrastive loss over $M$ image-text pairs:
$\mathcal{L}_{\text{pre-train}}=\frac{1}{M}\sum_{i=1}^{M}\Bigl{[}-\log\frac{\exp(s_{ii})}{\sum_{j=1}^{M}\exp(s_{ij})}\Bigr{]}\;+\;\frac{1}{M}\sum_{j=1}^{M}\Bigl{[}-\log\frac{\exp(s_{jj})}{\sum_{i=1}^{M}\exp(s_{ij})}\Bigr{]}\,.$ (2)
where $s_{ij}$ is the similarity score between $f(I_{i})$ and $g(T_{j})$, and The first and second terms represent the image-to-text and text-to-image contrastive losses ($\mathcal{L}_{\mathrm{img}\to\mathrm{text}}\ and\ \mathcal{L}_{\mathrm{text}\to\mathrm{img}}$). After pre-training, CLIP exhibits strong zero-shot performance by virtue of its large-scale domain knowledge and cross-modal alignment.
To enable rapid adaptation to downstream tasks, VLMs [56, 57] fine-tune a set of learnable prompts $P=[t_{1},t_{2},\dots,t_{b}]$, while keeping both the image and text encoders frozen. Given class name embeddings $[c_{1},c_{2},\dots,c_{N}]$, each class representation is formed by concatenating $P$ with $c_{i}$, and the prediction is computed as:
$p(y=i\mid I)=\frac{\exp\left(\cos\left(f(I),g([P,c_{i}])\right)/\tau\right)}{\sum_{j=1}^{N}\exp\left(\cos\left(f(I),g([P,c_{j}])\right)/\tau\right)},$ (3)
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes concatenation and $\tau$ is a temperature parameter. During fine-tuning, only $P$ is updated using cross-entropy loss, enabling task-specific adaptation.
3.3 Few-Shot SSL
@@
This section introduces four common semi-supervised settings and explains how they can be reformulated as a unified few-shot setting for pre-trained VLMs, referred to as Few-shot SSL. The dataset $\mathcal{D}$ used in this setting consists of a small labeled subset $\mathcal{D}_{l}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N_{l}}$ and a larger unlabeled subset $\mathcal{D}_{u}=\{\mathbf{x}_{j}\}_{j=1}^{N_{u}}$, where the number of labeled samples is significantly smaller than the unlabeled samples $(N_{l}\ll N_{u})$. The following provides a overview of each setting, and more details are shown in Appendix A.
@@
  • Standard Semi-Supervised Setting [48]. In this setting, the labels for both subsets belong to the same known class set $\mathcal{Y}^{l}$. For VLMs, which do not rely on unlabeled data for training, the Few-shots’ SSL framework reduces this setting to a standard few-shot scenario, provided the number of labeled samples remains the same.
@@
  • Open-Set Semi-Supervised Setting [41]. In this setting, the unlabeled dataset $\mathcal{D}_{u}$ may include instances from unseen classes, i.e., $\mathcal{Y}^{l}\subset\mathcal{Y}^{u}$ with $\mathcal{Y}^{\text{new}}=\mathcal{Y}^{u}\setminus\mathcal{Y}^{l}$. Since VLMs do not utilize unlabeled data, the open-set setting in Few-shots’ SSL reduces to the semi-supervised setting.
@@
  • Open-World Semi-Supervised Setting [39]. This setting builds upon the open-set setting, aiming to classify both known and new classes. For VLMs, pre-training endows them with zero-shot classification capability, allowing for direct classification of both known and new classes. Therefore, the Open-world semi-supervised setting is analogous to the commonly used Base-to-Novel setting [56] for VLMs.
@@
  • Long-Tailed Semi-Supervised Setting [29]. This setting features a highly imbalanced class distribution, where head classes have many samples and tail classes have few. The imbalance ratio (IR) is defined as the sample count ratio between the largest and smallest classes. Since VLMs use a fixed number of samples per class in few-shot settings, the head and tail classes become consistent due to the same number of samples. Thus, the Few-shot SSL framework reduces this setting to a standard few-shot scenario by matching the number of shots to that of the tail classes, ensuring fair comparison.
@@ 4.1 Baselines Selection
We selected FixMatch [48] (standard semi-supervised learning), DWD [3] (open-set semi-supervised learning), OwMatch [39] (open-world semi-supervised learning), and CCL [58] (long-tailed semi-supervised learning) as representative methods across different settings. For the pretrain–finetuning paradigm, we consider CLIP [43], a pre-trained VLM, along with its prompt tuning variants—CoOp [57], the first to introduce prompt tuning for CLIP, and PromptSRC [25], a state-of-the-art method. Including these methods in our comparison allows for a direct observation of the lower and upper bounds for pre-trained VLMs in semi-supervised settings, leading to a more comprehensive analysis.
4.2 SSL vs. VLM Finetuning in Semi-Supervised Setting
@@
$\Pi$-Model - 45.74 $\pm$ 3.97 85.99 $\pm$ 0.38 - 42.75 $\pm$ 0.48 62.12 $\pm$ 0.11 73.77 $\pm$ 0.82
Pseudo-Labeling - 50.22 $\pm$ 0.43 83.91 $\pm$ 0.28 - 42.62 $\pm$ 0.46 63.79 $\pm$ 0.19 72.01 $\pm$ 0.83
Mean Teacher - 67.68 $\pm$ 2.30 90.81 $\pm$ 0.19 - 46.09 $\pm$ 0.57 64.17 $\pm$ 0.24 78.57 $\pm$ 2.39
MixMatch 52.46 $\pm$ 11.50 88.95 $\pm$ 0.86 93.58 $\pm$ 0.10 32.39 $\pm$ 1.32 60.06 $\pm$ 0.37 71.69 $\pm$ 0.33 89.59 $\pm$ 0.61
UDA 70.95 $\pm$ 5.93 91.18 $\pm$ 1.08 95.12 $\pm$ 0.18 40.72 $\pm$ 0.88 66.87 $\pm$ 0.22 75.50 $\pm$ 0.25 92.34 $\pm$ 0.56
ReMixMatch 80.90 $\pm$ 9.64 94.56 $\pm$ 0.05 95.28 $\pm$ 0.13 55.72 $\pm$ 2.06 72.57 $\pm$ 0.31 76.97 $\pm$ 0.56 94.77 $\pm$ 0.45
FixMatch (RA) 86.19 $\pm$ 3.37 94.93 $\pm$ 0.65 95.74 $\pm$ 0.05 51.15 $\pm$ 1.75 71.71 $\pm$ 0.11 77.40 $\pm$ 0.12 92.02 $\pm$ 1.50
FixMatch (CTA) 88.61 $\pm$ 3.35 94.93 $\pm$ 0.33 95.69 $\pm$ 0.15 50.05 $\pm$ 3.01 71.36 $\pm$ 0.24 76.82 $\pm$ 0.11 94.83 $\pm$ 0.63
ZSCLIP 79.30 $\pm$ 0.00 46.00 $\pm$ 0.00 97.40 $\pm$ 0.00
CoOp 80.60 $\pm$ 1.79 85.10 $\pm$ 0.29 89.13 $\pm$ 0.12 49.00 $\pm$ 0.83 57.43 $\pm$ 0.25 61.40 $\pm$ 0.14 98.23 $\pm$ 0.17
PromptSRC 86.03 $\pm$ 0.48 90.10 $\pm$ 0.08 93.90 $\pm$ 0.14 60.33 $\pm$ 0.17 66.33 $\pm$ 0.29 70.10 $\pm$ 0.22 98.57 $\pm$ 0.12
Table 2: Performance of SSL and pre-trained VLMs on the ImageNet and STL-10 datasets.
(a) Performance on the ImageNet dataset.
Method Accuracy
UDA [53] 68.78 $\pm$ 0.43 (10% labels)
FixMatch [48] 71.46 $\pm$ 0.52 (10% labels)
ZSCLIP [43] 66.70 $\pm$ 0.00 (0 labels)
CoOp [57] 73.43 $\pm$ 0.12 (2% labels)
PromptSRC [25] 74.03 $\pm$ 0.09 (2% labels)
(b) Performance on STL-10 at different image resolutions.
Resolution Method
ZSCLIP CoOp PromptSRC
16 $\times$ 16 49.30 $\pm$ 0.00 64.83 $\pm$ 0.45 76.97 $\pm$ 0.45
32 $\times$ 32 83.40 $\pm$ 0.00 87.63 $\pm$ 0.17 92.30 $\pm$ 0.36
64 $\times$ 64 95.00 $\pm$ 0.00 96.73 $\pm$ 0.12 97.67 $\pm$ 0.05
96 $\times$ 96 97.40 $\pm$ 0.00 98.23 $\pm$ 0.17 98.57 $\pm$ 0.12
@@
Datasets and Experimental Setup. We evaluate both SSL and pre-trained VLMs on CIFAR-10/100 [27], STL-10 [9], and ImageNet [10], under varying annotation levels and augmentation strategies. For SSL methods (e.g., UDA [53], MixMatch [5]), we report the standard results as presented in FixMatch [48]. For VLMs, we simulate few-shot learning by training on the same labeled data only, without accessing unlabeled samples. All experiments on VLMs are conducted using the ViT-B/16 backbone and evaluated over three random seeds. Full implementation details are provided in Appendix A.2.
@@
Experiment Results. As shown in Table 1, SSL methods generally perform well on lower-resolution datasets such as CIFAR-10/100 (32 $\times$ 32), surpassing zero-shot CLIP and its fine-tuning variants. This advantage stems from the difficulty pre-trained VLMs face in extracting meaningful features from lower-resolution inputs. In contrast, pre-trained VLMs excels on higher-resolution datasets like STL-10 (96 $\times$ 96), where even zero-shot CLIP surpasses the best-performing SSL baselines. Moreover, as shown in Table 2(a), on the ImageNet dataset, FixMatch requires a significantly larger proportion of labeled data (around 10%) but still lags behind in accuracy. By comparison, VLMs achieve over 73% accuracy using only 2% of labeled samples, with performance improving as more labeled data becomes available. In addition to accuracy, SSL methods typically involve much longer training times. For instance, training SSL models on CIFAR datasets takes several hours [4, 48], whereas prompt tuning approaches for VLMs [24, 25, 57] can be completed in just a few minutes. These observations highlight the complementary strengths of SSL and pretrain-finetuning paradigms, with the latter generally offering a higher performance ceiling and faster training speed.
Analysis. To investigate how input resolution affects pre-trained models’ performance, we evaluate several methods on the STL-10 dataset at varying resolutions (see Table 2(b)). As resolution increases, all VLMs exhibit substantial accuracy gains, highlighting their sensitivity to image resolution. We also experiment with different visual encoder backbones (presented in Appendix B) and find that ViT-based architectures consistently outperform ResNet-50 across all resolutions, indicating that the performance degradation is not due to an intrinsic limitation of Vision Transformer [11]. In fact, it stems from the fixed input size requirement (e.g., 224 $\times$ 224) of VLMs: low-resolution images must be upsampled via interpolation to satisfy this requirement, a process that inevitably introduces distortions and impairs the models’ ability to extract reliable visual features.
@@
Takeaway 1: Pre-trained models demonstrate strong generalization capabilities and high data efficiency, enabling them to outperform traditional SSL methods with only a small amount of labeled data. However, achieving optimal performance with pre-trained VLMs requires higher-resolution input images, typically exceeding 96$\times$96.
4.3 SSL vs. VLM Finetuning in Open-Set Semi-Supervised Setting

### v4

# Unlabeled Data vs. Pre-trained Knowledge: Rethinking SSL in the Era of Large Models

@@ Abstract
Semi-supervised learning (SSL) alleviates the cost of data labeling process by exploiting unlabeled data and has achieved promising results. Meanwhile, with the development of large foundation models, exploiting pre-trained models becomes a promising way to address the label scarcity in the downstream tasks, such as various parameter-efficient fine-tuning techniques. This raises a natural yet critical question: When labeled data is limited, should we rely on unlabeled data or pre-trained models? To investigate this issue, we conduct a fair comparison between SSL methods and pre-trained models (e.g., CLIP) on representative image classification tasks under a controlled supervision budget. Experiments reveal that SSL has met its “Waterloo” in the era of large models, as pre-trained models show both high efficiency and strong performance on widely adopted SSL benchmarks. This underscores the urgent need for SSL researchers to explore new avenues, such as deeper integration between the SSL and pre-trained models. Furthermore, we investigate the potential of Multi-Modal Large Language Models (MLLMs) in image classification tasks. Results show that, despite their massive parameter scales, MLLMs still face significant performance limitations, highlighting that even a seemingly well-studied task remains highly challenging.
@@ 1 Introduction
Deep neural networks have demonstrated performance comparable to that of humans in certain supervised learning tasks, such as image classification (Geirhos et al., 2018; He et al., 2016; Lee et al., 2013b). However, these impressive results rely heavily on large amounts of labeled training data. In many real-world applications, obtaining such labeled data often demands substantial human and financial resources (Yang et al., 2022). To reduce the dependence on labeled data and thereby cut costs, researchers have proposed Semi-Supervised Learning (SSL), leveraging unlabeled data to enhance model performance for image classification tasks.
In recent years, Pre-trained Models have attracted considerable attention. Among them, vision-language models (VLMs) such as CLIP (Radford et al., 2021), which achieves semantic alignment during pre-training, has become a widely adopted backbone for image classification tasks. Building upon CLIP, various fine-tuning strategies (e.g., prompt tuning (Zhou et al., 2022b)) have been proposed to enhance task-specific performance using minimal labeled data by leveraging the rich knowledge acquired during pre-training (Zhou et al., 2022b; Khattak et al., 2023a; b). More recently, multi-modal large language models (MLLMs) such as Qwen-VL (Bai et al., 2023) and GPT-4o-min (Hurst et al., 2024) have emerged, featuring significantly larger parameter scales and stronger general capabilities in visual tasks.
@@
As shown in Figure 1, both SSL and pre-trained models are well suited to image classification tasks with limited labeled data, by leveraging unlabeled data and pre-trained knowledge, respectively. This makes us come up with three questions naturally:
  1. 1.
    When labels are scarce, should we prioritize utilizing unlabeled data or pre-trained models?
  2. 2.
    What are the strengths and limitations of leveraging unlabeled data versus exploiting pre-trained models?
  3. 3.
    How can we design methods that mitigate the shortcomings of both approaches while maximizing their complementary benefits?
To answer these questions, we adopt widely used CLIP-based VLMs as representative pre-trained models, owing to their strong generalization capabilities and proven effectiveness in zero- and few-shot learning across diverse domains (Radford et al., 2021; Zhou et al., 2022b; Khattak et al., 2023b). We extracted and designed a fair evaluation framework based on papers from top-tier conferences and journals to compare various SSL methods and pre-trained models under consistent experimental conditions. Specifically, we standardize the amount of labeled data used across different experimental setups, including semi-supervised, open-set semi-supervised, open-world semi-supervised, and long-tailed semi-supervised settings, ensuring equal supervision signals for all approaches.
The experimental results show that the SSL method faces its Waterloo, with problems such as low training efficiency and poor performance. Moreover, when the test data contains unseen classes, SSL struggles to learn and adapt effectively. In contrast, pre-trained models, endowed with rich prior knowledge, can quickly adapt to downstream tasks with limited data, demonstrating strong generalization to open-world scenarios and long-tailed distributions. Although pre-trained models may suffer from adaptation bias due to distribution discrepancies between pre-training data and downstream tasks, these issues can be alleviated through additional training or data distillation techniques. Such approaches also yield performance gains and higher efficiency compared to SSL methods.
To bridge the gap between the two paradigms, we further explore hybrid algorithms. By leveraging their generalization ability, these methods can generate more reliable pseudo-labels, thereby substantially enhancing the efficiency and stability of semi-supervised training. We explore the performance of larger-scale MLLMs on image classification tasks, and the results show that they are likewise constrained by adaptation bias and other issues arising from the limitations of their visual encoders. To further advance research in this area, we propose the following directions:
  1. 1.
    It is no longer sufficient for SSL researchers to compare only with previous SSL methods in image classification; systematic comparisons with pre-trained models are necessary to ensure fair and meaningful evaluations.
  2. 2.
    Exploring effective strategies to integrate SSL techniques with the adaptation of pre-trained models represents another promising approach to enhance both the efficiency of SSL and the adaptability of pre-trained models.
  3. 3.
    Exploring the capabilities of MLLMs on image classification tasks, particularly across varying distributions and resolutions, as experiments highlight that even a seemingly well-studied task like image classification remains challenging for MLLMs.
@@ 2 Related Work
Semi-Supervised Learning. The goal of Semi-Supervised Learning is to reduce the cost of data annotation by leveraging unlabeled data, under the assumption that both labeled and unlabeled samples belong to a predefined set of known classes. The primary objective is to assign accurate ground-truth labels to unlabeled instances. Existing SSL methods can be broadly categorized into three main approaches: Hybrid and Holistic Methods (Berthelot et al., 2019a; b). Pseudo-Labeling assigns pseudo-labels to unlabeled data based on the model’s own predictions throughout training (Lee et al., 2013a). FixMatch (Sohn et al., 2020) combines regularization and confidence-based filtering by enforcing consistency between weakly and strongly augmented samples, while only using high-confidence pseudo-labels for training, resulting in significant performance gains with limited labeled data.
Pre-trained Models. Recent advances in pre-trained models have attracted considerable attention, particularly with the emergence of VLMs such as CLIP (Radford et al., 2021) and MLLMs such as Qwen-VL (Bai et al., 2023). While prior work (Zoph et al., 2020) has compared pre-trained models with self-training algorithms in object detection and segmentation tasks, it did not consider VLMs, MLLMs, or even SSL settings. To address this gap, we focus on the widely studied image classification task and systematically investigate the performance of VLMs and their fine-tuning strategies. As the pioneering work for fine-tuning VLMs, CoOp (Zhou et al., 2022b) for the first time introduces learnable prompts to transfer the task-specific knowledge to VLMs. Building on this idea, MaPLe (Khattak et al., 2023a) and PromptSRC (Khattak et al., 2023b) extend prompt tuning to both visual and textual encoders, enabling joint vision-language adaptation. Multi-modal large language models, on the other hand, are primarily geared toward reasoning tasks, and their capabilities in image classification remain underexplored. Therefore, this paper selects three MLLMs for a comparative study.
3 Evaluation Settings
@@
We summarize four common SSL settings from 50 recent top-tier publications, which represent the primary and most challenging focuses of current research. Then, we reformulate these settings for pre-trained VLMs to ensure fair comparisons. The dataset $\mathcal{D}$ used in this setting consists of a small labeled subset $\mathcal{D}_{l}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N_{l}}$ and a larger unlabeled subset $\mathcal{D}_{u}=\{\mathbf{x}_{j}\}_{j=1}^{N_{u}}$, where the number of labeled samples is significantly smaller than the unlabeled samples $(N_{l}\ll N_{u})$. The following provides an overview of each setting, and details are shown in Appendix B.
@@
  • Standard Semi-Supervised Learning (Sohn et al., 2020). In this setting, the labels for both subsets belong to the same known class set $\mathcal{Y}^{l}$. For pre-trained models, which do not rely on unlabeled data for training, our setting reduces this setting to a standard few-shot scenario, provided the number of labeled samples remains the same.
@@
  • Open-Set Semi-Supervised Learning (Oliver et al., 2018). In this setting, the unlabeled dataset $\mathcal{D}_{u}$ may include instances from unseen classes, i.e., $\mathcal{Y}^{l}\subset\mathcal{Y}^{u}$ with $\mathcal{Y}^{\text{new}}=\mathcal{Y}^{u}\setminus\mathcal{Y}^{l}$. Since pre-trained models do not utilize unlabeled data, our open-set setting reduces to the semi-supervised setting.
@@
  • Open-World Semi-Supervised Learning (Niu et al., 2024). This setting extends the open-set scenario by aiming to classify both known and novel classes. Pre-trained models naturally support this via zero-shot classification, making the open-world semi-supervised setting analogous to the widely used base-to-novel setting (Zhou et al., 2022a).
@@
  • Long-Tailed Semi-Supervised Learning (Kukleva et al., 2023). This setting features a highly imbalanced class distribution, where major (head) classes have many samples and minor (tail) classes have few. The imbalance ratio (IR) is defined as the sample count ratio between the largest and smallest classes. Since VLMs use a fixed number of samples per class in few-shot settings, aligning the shot count with tail classes reduces the problem to a standard few-shot scenario, enabling fair comparison.
@@ 4.1 Baselines Selection
Given that SSL is primarily applied to image classification tasks, we selected FixMatch (Sohn et al., 2020) as a representative method for standard semi-supervised learning, DWD (Ban et al., 2024) for open-set semi-supervised learning, OwMatch (Niu et al., 2024) for open-world semi-supervised learning, and CCL (Zhou et al., 2024) for long-tailed semi-supervised learning. For pre-trained models, we consider CLIP (Radford et al., 2021) along with its prompt tuning variants: CoOp, which was the first to introduce prompt tuning for CLIP (Zhou et al., 2022b), and PromptSRC (Khattak et al., 2023b), a state-of-the-art method in this line of work.
4.2 Comparisons in Standard SSL Setting
@@
MixMatch 52.46±11.50 88.95±0.86 93.58±0.10 32.39±1.32 60.06±0.37 71.69±0.33 89.59±0.61
UDA 70.95±5.93 91.18±1.08 95.12±0.18 40.72±0.88 66.87±0.22 75.50±0.25 92.34±0.56
ReMixMatch 80.90±9.64 94.56±0.05 95.28±0.13 55.72±2.06 72.57±0.31 76.97±0.56 94.77±0.45
FixMatch (RA) 86.19±3.37 94.93±0.65 95.74±0.05 51.15±1.75 71.71±0.11 77.40±0.12 92.02±1.50
FixMatch (CTA) 88.61±3.35 94.93±0.33 95.69±0.15 50.05±3.01 71.36±0.24 76.82±0.11 94.83±0.63
ZSCLIP 79.30±0.00 46.00±0.00 97.40±0.00
CoOp 80.60±1.79 85.10±0.29 89.13±0.12 49.00±0.83 57.43±0.25 61.40±0.14 98.23±0.17
PromptSRC 86.03±0.48 90.10±0.08 93.90±0.14 60.33±0.17 66.33±0.29 70.10±0.22 98.57±0.12
Table 2: Performance of SSL and pre-trained VLMs on the ImageNet dataset.
Method UDA FixMatch ZSCLIP CoOp PromptSRC
Label 10% labels 10% labels 0 labels 2% labels 2% labels
Accuracy 68.78±0.43 71.46±0.52 66.70±0.00 73.43±0.12 74.03±0.09
@@
Datasets and Experimental Setup. We evaluate both SSL and pre-trained VLMs on CIFAR-10/100 (Krizhevsky et al., 2009), STL-10 (Coates et al., 2011), and ImageNet (Deng et al., 2009), under varying annotation levels and augmentation strategies. For VLMs, we simulate few-shot learning by training on the same labeled data only, without accessing unlabeled samples. Full implementation details are provided in Appendix B.2.
@@
Experiment Results. As shown in Table 1, SSL methods generally perform well on low-resolution datasets such as CIFAR-10 and CIFAR-100, which have an image resolution of 32 $\times$ 32, surpassing zero-shot CLIP and its fine-tuning variants. This advantage stems from the challenge of extracting meaningful features from low-resolution inputs due to the “adaptation bias” between the pre-training data of VLMs and the target downstream tasks. In contrast, pre-trained VLMs excel on higher-resolution datasets like STL-10 (96 $\times$ 96), where even zero-shot CLIP surpasses the best-performing SSL baselines. Moreover, as shown in Table 2, on the ImageNet dataset, FixMatch requires a significantly larger proportion of labeled data (around 10%) but still lags behind in accuracy. By comparison, VLMs achieve over 73% accuracy using only 2% of labeled samples, with performance improving as more labeled data becomes available. In addition to accuracy, training SSL models on CIFAR datasets takes GPU hours (Berthelot et al., 2019a; Sohn et al., 2020), whereas prompt tuning approaches for VLMs (Khattak et al., 2023a; b; Zhou et al., 2022b) can be completed in just a few GPU minutes. In summary, the above observations highlight the complementary strengths of SSL and pre-trained models, with the latter generally offering a higher performance ceiling and faster training speed.
Figure 2: Performance of different backbones.
Figure 3: Performance at different resolutions.
Analysis of “Adaptation Bias”. We argue that “adaptation bias” is typically caused by two main factors: the capacity of the visual encoder and discrepancies between training and testing styles (e.g., differences in resolution and image style). To evaluate the impact of resolution, we test CLIP models with different backbones on the CIFAR-10/100 datasets. As shown in Table 3, performance improves steadily with the number of visual encoder parameters, indicating that the suboptimal results on low-resolution images are not due to the encoder itself. Since the original training data of CLIP is unavailable, we approximate this bias through input resolution (CLIP is fixed at an input resolution of 224 $\times$ 224). We then evaluated various methods under different input resolutions on the STL-10 dataset (Figure 3). As the resolution increases, model performance consistently improves, and adaptation bias is gradually reduced. This suggests that the distribution shift between training and testing is the fundamental cause of “adaptation bias”. Fine-tuning can partially mitigate the adaptation bias of models toward low-resolution images (e.g., PromptSRC outperforms CLIP at 16 $\times$ 16 resolution). Additionally, model distillation techniques have been shown to alleviate this bias (Lee et al., 2025; Zhou et al., 2023). For instance, RADIO (Heinrich et al., 2024) enhances the stability and performance of student models across varying input resolutions by distilling knowledge from teacher models and incorporating strategies such as multi-resolution training and balanced teacher losses.
@@
Takeaway 1: Pre-trained models demonstrate strong generalization capabilities and high data efficiency, enabling them to outperform traditional SSL methods with only a small amount of labeled data. Despite the presence of “adaptation bias” in pre-trained models, downstream adaptation techniques can yield performance gains and improve efficiency.
4.3 Comparisons in Open-Set SSL Setting